# Unveiling the Basin-Like Loss Landscape in Large Language Models

### v2

@@ 4.1 Any Alignment Can Be Bounded by Average-case Alignment
Examples. Through hypothesis testing in Sec.˜D.4, we establish that Qwen2.5-7B (Yang et al., 2024) has a $\sigma$-basin with $\sigma=0.003$ on the safety task using the AdvBench dataset, where $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\geq 0.9$. Based on this, we derive a lower bound on safety degradation as a function of $\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}$. As shown in Fig.˜4(a), for a $\sigma=0.003$ basin, a higher performance on the original parameters, i.e., $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]$, yields a stronger guarantee on the fine-tuned parameters $\bm{\theta}_{sft}$, i.e., $J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$.
@@
Increasing the Basin Size Improves the Guarantee. As illustrated in Fig.˜4(b), enlarging the basin of a model while preserving the performance on the original model, i.e., $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$, linearly strengthens the theoretical guarantee. In other words, increasing the basin size $\sigma$ results in a linear improvement in the performance degradation guarantee during fine-tuning.
@@
Figure 4: Lower bound guarantees. (a) The lower bound on the benchmark value of the smoothed fine-tuned model $J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$ for varying benchmark values on the smoothed original model $\bm{\theta}_{0}$, i.e., $p_{A}:=\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]$, with $\sigma=0.003$. (b) The lower bound on the benchmark value of the smoothed fine-tuned model $J_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$ for varying basin sizes $\sigma$, with $p_{A}=0.9$. (c) Histogram of L2 distances between token embeddings.
4.2 Input-space Robustness Can Be Bounded by Most-case Alignment
In this section, we propose that the size of a basin also provides a (loose) guarantee of performance degradation against jailbreaking attacks (Zou et al., 2023; Wei et al., 2023b), by considering only the very first layer of the model.
Intuition. Let $\bm{W}$ denote the embedding layers. Given that the embedding layers of current large language models are onto transformations (i.e., $\bm{W}$ is column full-rank) (Carlini et al., 2024), the activation after a weight perturbation, $\bm{W}\bm{x}+\bm{\delta_{W}}\bm{x}$, and the activation after an input perturbation, $\bm{W}\bm{x}+\bm{W}\bm{\delta_{x}}$, can yield the same vector (Zhang et al., 2024a). Thus, if the model is robust to any weight perturbation $\|\bm{\delta_{W}}\|_{2}\leq\epsilon_{\bm{W}}$, it is also robust to any input perturbation $\bm{\delta_{x}}$ such that $\bm{W}\bm{\delta_{x}}\in\{\bm{\delta_{W}}\bm{x}\mid\|\bm{\delta_{W}}\|_{2}\leq\epsilon_{\bm{W}}\}$.
Theorem 4.5
Let $\mathcal{D}^{\prime}$ be a modified version of $\mathcal{D}$ where $k$ tokens are substituted, i.e., each token $\bm{e}_{i}$ is replaced with $\bm{e}_{i}^{\prime}$ in the set $\mathcal{C}=\{(\bm{e}_{i},\bm{e}_{i}^{\prime})\}_{i=1}^{k}$. For any benchmark $J_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, the performance degradation can be bounded by the embedding differences:
$\mathbb{E}_{\bm{\epsilon}}[J_{f,\mathcal{D}^{\prime}}(\bm{\theta}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)-\frac{\sqrt{\sum_{i=1}^{k}\|\bm{W}\bm{e}_{i}-\bm{W}\bm{e}_{i}^{\prime}\|_{2}^{2}}}{\sigma}\right).$ (7)
@@ 4.2 Input-space Robustness Can Be Bounded by Most-case Alignment
Application to Certification Against Jailbreaking. Consider $\mathcal{D}=\{\bm{x}\}$ containing a single input instance and $J$ as a safety detector that returns values greater than 0.5 for safe outputs and less than 0.5 for harmful outputs. We can certify robustness against jailbreaking attacks by determining whether $J_{f,\mathcal{D}^{\prime}}(\bm{\theta}+\bm{\epsilon})$ remains above 0.5 after modifying $\mathcal{D}=\{\bm{x}\}$ to $\mathcal{D}^{\prime}=\{\bm{x}_{adv}\}$.
We derive the following conclusions for Qwen2.5-7B (Yang et al., 2024):
@@
Substituting Some Tokens Preserves Performance. Most large language models, including Qwen2.5-7B, use BPE tokenizers (Sennrich, 2015), where tokens with and without leading spaces are distinct (e.g., “hi” vs. “_hi”). These tokens often have small $\ell_{2}$ distances between each other. Special tokens, such as “<s>”, “.”, “”, “ ”, and “ ”, also have small $\ell_{2}$ distances. Substituting these tokens, as evaluated in Eq.˜7, results in equivalent parameter perturbations too small to significantly alter model outputs. Thus, the model exhibits robustness to these token pairs.
Current LLMs Are Generally Sensitive to Input Changes. As shown in Fig.˜4(c), only a small fraction of token substitutions have negligible effects on model outputs. This flexibility allows models to generate diverse responses based on input variations but also introduces robustness vulnerabilities. Minor input changes can lead to significant output variations, potentially resulting in incorrect or unsafe responses (Zou et al., 2023; Chen et al., 2025a; Andriushchenko et al., 2024).
@@
(c) SFT Loss Curve
Figure 5: Pre-training and fine-tuning using GO and Adam optimizers. (a) The pre-training loss curve on OpenWebText. (b) The loss landscape on OpenWebText after pre-training. (c) Performance changes in old capability (OpenWebText) and new capability (Alpaca).
4.3 Basin May Have Sufficient Expressive Power in the Future
Given the theoretical guarantees of fine-tuning within a basin, one might wonder whether subsequent fine-tuning can be constrained to a theoretically guaranteed region, thereby preserving all original capabilities and achieving continual learning without forgetting (Chen et al., 2025b). A primary concern with this constraint method is whether it limits the expressive power of the hypothesis set. Specifically, if the hypothesis set is constrained to functions where $\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}\leq O(\sigma)$, can it still effectively fit the fine-tuning dataset?
We propose that the larger the model and its basin, the greater the expressive power within the basin, enabling better acquisition of new capabilities. As demonstrated in the following lemmas, a large model with a large basin may be even more expressive than a smaller model. Thus, in the future, one may prefer training a significantly larger model and enlarging its basin, either actively or passively (as larger models naturally form larger basins). In this case, constraining fine-tuning within the basin will provide sufficient expressive power to acquire new capabilities.
Lemma 4.6
(Bartlett et al., 2019) The upper and lower bounds of the VC dimension of an $l$-layer fully connected ReLU network are $O(dl\log d)$ and $\Omega(dl\log\frac{d}{l})$, respectively.
Lemma 4.7
(Adapted from Neyshabur et al. (2015)) The upper bound of the Rademacher complexity of a two-layer ReLU neural network, where each parameter has an absolute value bound of $O(\sigma)$, is $O(d\sigma^{2})$.
5 Enlarging Basins
Given that a larger basin provides stronger robustness guarantees against any fine-tuning and jailbreaking attacks, one might wonder whether the basin can be enlarged. In this section, we discover that the basin can be readily expanded, may due to the over-parameterization property of current LLMs (Liu et al., 2022; Allen-Zhu et al., 2019).
@@ 5.1 Gaussian-augmented Optimizer
As outlined in Definition 3, enlarging the basin size requires optimizing $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[J_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$, ensuring that the model $\bm{\theta}$ is robust to Gaussian perturbations. To this end, we define the loss function as the expected cross-entropy loss over perturbed parameters:
@@
As detailed in Algorithm 1, the optimization process involves performing a forward pass with perturbed parameters $\bm{\theta}+\bm{\epsilon}$, computing the loss, calculating the gradient via backpropagation, and using the gradient to update the parameters with a standard optimizer. We name this process GO optimizer.
@@ 5.2 Experimental Results
Settings. To better study the effect of the GO optimizer, we applied it during pre-training. Following the NanoGPT pipeline (Karpathy, 2023), we pre-train a GPT2-127M model on OpenWebText for 600,000 steps using the same hyperparameters as in the repository. For the GO optimizer, we use identical hyperparameters, except for the additional hyperparameter $\sigma=0.01$. We then fine-tune these two models on the Alpaca dataset to enhance conversational capabilities while examining performance degradation, both using the Adam optimizer.
Pre-training Results. As shown in Fig.˜5(b), adding Gaussian noise to parameters can significantly expand the basin size. Although the GO optimizer is slower than Adam at the beginning of training, as it requires optimizing the loss across the entire neighborhood, it gradually catches up and reduces the performance gap (see Fig.˜5(a)), possibly due to the over-parameterization property of current large language models. More interestingly, the GO optimizer appears to introduce some inductive bias, outperforming AdamW on some benchmarks (see Sec.˜D.5).
Fine-tuning Results. We propose that the larger the basin, the less performance degradation occurs during benign fine-tuning (since benign fine-tuning, which aims to enhance the model, should not be more harmful than Gaussian noise) and that it does not affect the acquisition of new capabilities (since larger basins imply greater expressive power within the basin). As demonstrated in Fig.˜5(c), fine-tuning a model pre-trained with the Gaussian-augmented optimizer achieves significantly less performance degradation on OpenWebText while maintaining nearly the same learning speed on Alpaca, supporting this claim. See Sec.˜D.5 for detailed analysis of additional benchmarks.

### v3

@@ 4.1 Any Alignment Can Be Bounded by Average-case Alignment
Examples. Through hypothesis testing in Sec.˜D.4, we establish that Qwen2.5-7B (Yang et al., 2024) has a $\sigma$-basin with $\sigma=0.003$ on the safety task using the AdvBench dataset, where $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\geq 0.9$. Based on this, we derive a lower bound on safety degradation as a function of $\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}$. As shown in Fig.˜4(a), for a $\sigma=0.003$ basin, a higher performance on the original parameters, i.e., $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]$, yields a stronger guarantee on the fine-tuned parameters $\bm{\theta}_{sft}$, i.e., $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$.
@@
Increasing the Basin Size Improves the Guarantee. As illustrated in Fig.˜4(b), enlarging the basin of a model while preserving the performance on the original model, i.e., $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$, linearly strengthens the theoretical guarantee. In other words, increasing the basin size $\sigma$ results in a linear improvement in the performance degradation guarantee during fine-tuning.
@@
Figure 4: Lower bound guarantees. (a) The lower bound on the benchmark value of the smoothed fine-tuned model $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$ for varying benchmark values on the smoothed original model $\bm{\theta}_{0}$, i.e., $p_{A}:=\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]$, with $\sigma=0.003$. (b) The lower bound on the benchmark value of the smoothed fine-tuned model $\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})$ for varying basin sizes $\sigma$, with $p_{A}=0.9$. (c) Histogram of L2 distances between token embeddings.
4.2 Input-space Robustness: A Heuristic Analysis via Embedding Geometry
In this section, we explore the theoretical connection between parameter-space basins and input-space robustness. While full certification in the discrete token space is intractable, we provide a heuristic analysis based on the local geometry of the embedding layer (Zou et al., 2023; Wei et al., 2023b), considering only the very first layer of the model.
Intuition. Let $\bm{W}$ denote the embedding layers. Given that the embedding layers of current large language models are onto transformations (i.e., $\bm{W}$ is column full-rank) (Carlini et al., 2024), the activation after a weight perturbation, $\bm{W}\bm{x}+\bm{\delta_{W}}\bm{x}$, and the activation after an input perturbation, $\bm{W}\bm{x}+\bm{W}\bm{\delta_{x}}$, can yield the same vector (Zhang et al., 2024a). Thus, if the model is robust to any weight perturbation $\|\bm{\delta_{W}}\|_{2}\leq\tau_{\bm{W}}$, it is also locally robust to any input perturbation $\bm{\delta_{x}}$ such that $\bm{W}\bm{\delta_{x}}\in\{\bm{\delta_{W}}\bm{x}\mid\|\bm{\delta_{W}}\|_{2}\leq\tau_{\bm{W}}\}$.
Theorem 4.5.
Let $\mathcal{D}^{\prime}$ be a modified version of $\mathcal{D}$ where $k$ tokens are substituted, i.e., each token $\bm{e}_{i}$ is replaced with $\bm{e}_{i}^{\prime}$ in the set $\mathcal{C}=\{(\bm{e}_{i},\bm{e}_{i}^{\prime})\}_{i=1}^{k}$. For any benchmark $\mathcal{S}_{\mathcal{D}}:\mathbb{R}^{d}\to[0,1]$, the performance degradation under local embedding perturbations can be approximated by:
$\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}^{\prime}}(\bm{\theta}+\bm{\epsilon})]\geq\Phi\left(\Phi^{-1}\left(\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]\right)-\frac{\sqrt{\sum_{i=1}^{k}\|\bm{W}\bm{e}_{i}-\bm{W}\bm{e}_{i}^{\prime}\|_{2}^{2}}}{\sigma}\right).$ (6)
@@ 4.2 Input-space Robustness: A Heuristic Analysis via Embedding Geometry
Application to Certification Against Jailbreaking. Consider $\mathcal{D}=\{\bm{x}\}$ containing a single input instance and $\mathcal{S}$ as a safety detector that returns values greater than 0.5 for safe outputs and less than 0.5 for harmful outputs. We can analyze robustness against jailbreaking attacks by determining whether $\mathcal{S}_{f,\mathcal{D}^{\prime}}(\bm{\theta}+\bm{\epsilon})$ remains above 0.5 after modifying $\mathcal{D}=\{\bm{x}\}$ to $\mathcal{D}^{\prime}=\{\bm{x}_{adv}\}$.
@@
Substituting Some Tokens Preserves Performance. Most LLMs, including Qwen2.5-7B, use BPE tokenizers (Sennrich, 2015), where tokens with and without leading spaces are distinct (e.g., “hi” vs. “_hi”). These tokens often have small $\ell_{2}$ distances between each other. Special tokens, such as “<s>”, “.”, “”, “ ”, and “ ”, also have small $\ell_{2}$ distances. Substituting these tokens, as evaluated in Eq.˜6, results in equivalent parameter perturbations too small to significantly alter model outputs. Thus, the model exhibits robustness to these token pairs. However, only a small fraction of token substitutions have such negligible effects on model outputs. This allows models to generate diverse responses based on input variations while introducing robustness vulnerabilities.
Limitations on Tokenization Attacks.
We explicitly acknowledge that our theoretical bound relies on the local flatness within the embedding space. Thus, it primarily certifies robustness against perturbations where the tokenization structure remains relatively stable (e.g., substitution of semantically similar tokens). It does not extend to global or “off-manifold” perturbations that drastically disrupt the tokenization itself—such as attacks that greatly alter the token count or leverage random capitalization to trigger different tokenization boundaries (e.g., Hughes et al. (2024)). In such cases, the resulting shift in the discrete token space will exceed the certified radius in the embedding space. Thus, Eq.˜6 serves as a geometric intuition regarding local stability.
4.3 Scaling and Expressive Power within the Basin
Given the theoretical guarantees of fine-tuning within a basin, a critical question arises: can subsequent fine-tuning be constrained to this theoretically guaranteed region to achieve continual learning without forgetting (Chen et al., 2025b)? A primary concern is whether such a constraint limits the expressive power of the hypothesis set. Specifically, if the hypothesis set is restricted to functions where $\|\bm{\theta}_{sft}-\bm{\theta}_{0}\|_{2}\leq O(\sigma)$, can it still effectively fit the fine-tuning dataset?
We propose that the larger the model and its basin, the greater the expressive power within the basin, enabling better acquisition of new capabilities. Foundational results in learning theory (Bartlett et al., 2019; Neyshabur et al., 2015; Kajitsuka & Sato, 2025) establish that the expressive capacity of a neural network scales with both its parameter dimension and the allowable weight norm (which corresponds to our basin radius $\sigma$). Therefore, we argue that a significantly larger model operating within a correspondingly larger basin retains a hypothesis space rich enough to acquire new capabilities.
Thus, in the future, one may prefer training larger models and enlarging their basins, either actively (via GO, see Sec.˜5) or passively (as larger models naturally form larger basins, see Sec.˜5.3). In this regime, constraining fine-tuning within the basin offers a viable path to preserving prior capabilities without sacrificing learnability.
5 Enlarging Basins
In Sec.˜4, we decompose the expected performance degradation during SFT as:
$\underbrace{\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0})-\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]}_{\text{Total Degradation}}=\underbrace{\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]-\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{sft}+\bm{\epsilon})]}_{\text{Bounded by Theorem 4.3}}+\underbrace{\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0})-\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]}_{\text{Resilience to Gaussian Noise}}$
Crucially, when $\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]\to 1$, randomized smoothing theory guarantees that the first term vanishes (see Eq.˜5), while the second term naturally approaches zero. Consequently, optimizing $\mathbb{E}_{\bm{\epsilon}}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}_{0}+\bm{\epsilon})]$ (enlarging the basin) theoretically mitigates both sources of degradation.
Guided by this insight, in this section, we empirically validate this hypothesis. We discover that: (1) the basin can be readily expanded, potentially due to the over-parameterization of current LLMs (Allen-Zhu et al., 2019; Liu et al., 2022); and (2) explicitly optimizing for robustness against Gaussian perturbations effectively mitigates catastrophic forgetting—lower performance degradation under Gaussian noise directly translates to reduced forgetting under fine-tuning.
@@ 5 Enlarging Basins
(c) SFT Degradation
Figure 5: Comparison of GO and other sharpness-aware optimizers (SAM-$\rho$, CDrop-$\sigma$). Suffixes denote the hyperparameter magnitude. (a) Pre-training loss curve. (b, c) Validation of the upper bound hypothesis: Resilience to Gaussian noise directly serves as an empirical upper bound for degradation during subsequent fine-tuning. Raw data are presented in Table˜4.
@@ 5.1 Gaussian-augmented Optimizer
As outlined in Definition 3, to effectively enlarge the basin size—or equivalently, enhance resilience to Gaussian noise—the most direct approach is to optimize $\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}[\mathcal{S}_{f,\mathcal{D}}(\bm{\theta}+\bm{\epsilon})]$, ensuring that the model $\bm{\theta}$ is robust to Gaussian perturbations. To this end, we define the loss function as the expected cross-entropy loss over perturbed parameters:
@@
This results in Algorithm 1, which involves performing a forward pass with perturbed parameters $\bm{\theta}+\bm{\epsilon}$, computing the loss, calculating the gradient via backpropagation, and using the gradient to update the parameters with a standard optimizer. We name this process GO optimizer.
@@ 5.2 Experimental Results
Settings. To better study the effect of the GO optimizer, we applied it during pre-training. Following the NanoGPT pipeline (Karpathy, 2023), we pre-train a GPT2-127M model on OpenWebText for $8\times$ Chinchilla steps using the same hyperparameters as in the repository. For the GO optimizer, we use identical hyperparameters, except for $\sigma=0.01$. We then fine-tune these models on the Alpaca dataset to enhance conversational capabilities while examining performance degradation, both using the Adam optimizer. For comparison, we also include Sharpness-Aware Minimization (Foret et al., 2020) and Continuous Dropout (Srivastava et al., 2014) as baselines.
GO Optimizer Significantly Expands Basin Size. As shown in Fig.˜5(b), adding Gaussian noise to parameters can significantly expand the basin size. Although the GO optimizer is slower than Adam at the beginning of training—as it requires optimizing the loss across the entire neighborhood—it gradually catches up and reduces the performance gap (see Fig.˜5(a)), possibly due to the over-parameterization property of current LLMs. Furthermore, the GO optimizer appears to introduce a beneficial inductive bias, outperforming AdamW on some benchmarks (see Sec.˜D.5).
Gaussian Resilience Bounds Fine-tuning Degradation. We posit that the performance degradation induced by random Gaussian noise serves as an empirical upper bound for benign fine-tuning degradation. The rationale is that since benign SFT aims to enhance the model, it should theoretically be “less harmful” to existing capabilities than blind random noise. Consequently, explicitly minimizing Gaussian-induced degradation should constrain the degradation observed during downstream fine-tuning. Crucially, as shown in Figs.˜5(b), 5(c) and 4, there is a strict correspondence where strictly suppressing degradation under Gaussian noise directly translates to minimized performance degradation during subsequent fine-tuning.
Comparison with Other Sharpness-Aware Optimizers. Unlike methods that target worst-case sharpness (Foret et al., 2020) or act as implicit proxies (Srivastava et al., 2014), GO explicitly optimizes for average-case resilience against Gaussian perturbations. Given our central premise that this average-case Gaussian resilience serves as an upper bound for benign fine-tuning degradation, GO aligns most directly with the objective of preserving capabilities. Consequently, it demonstrates superior efficacy in mitigating catastrophic forgetting compared to other landscape-aware optimizers, as evidenced by the lowest SFT degradation in Fig.˜5(c). See Sec.˜D.8 for details.
5.3 Discussions
Basin Evolution During Training. We investigate the temporal dynamics of basin formation throughout the pre-training trajectory. As visualized in Fig.˜7, the basin is not a static property determined at initialization; rather, it is an emergent structure that gradually widens as training progresses. This continuous expansion aligns with the theoretical understanding that the stochastic noise in gradient descent algorithms introduces an implicit bias towards flatter minima (Damian et al., 2021; Li et al., 2021; 2025b). This observation suggests a potential benefit of "over-training": while the loss improvement may saturate, the geometric properties of the landscape (i.e., basin width) may continue to improve, thereby enhancing the model’s robustness to future fine-tuning degradation.
Fine-tuning with Different Distribution Gaps. As demonstrated in Table˜4, when fine-tuning on different datasets, the larger the distribution gap, the sharper the forgetting, aligning with our analysis in Sec.˜3.4. However, suppressing degradation under Gaussian noise consistently translates to resistance against subsequent fine-tuning, regardless of the fine-tuning dataset.